Find m: 2x/m + 3a/x = 5/c
Answer by stanbon(75887)

Find m: 2x/m + 3a/x = 5/c
-----
2x/m = (5/c)-(3a/x)
---
2x/m = (5x-3ac)/(xc)
---
m/(2x) = (xc)/(5x-3ac)
---
m = (2x^2c)/(5x-3ac)
============================
Cheers,
Stan H.
================